# Twisted Spectra Revisited

### v1

@@ 6.2. Dualizability of Twisted Spectra
[…]
@@
$\Psi_{X}\circ\Phi_{Y}$ $*\displaystyle\simeq A_{!}\circ(-\otimes Y)\circ(-\otimes X)\circ A^{*}$
$*\displaystyle\simeq(-\otimes A_{!}(X\otimes Y))\circ A^{*}$
@@
$\Phi_{X\boxtimes Y}$ $*\displaystyle=(p_{2})_{!}\circ(-\otimes(X\boxtimes Y))\circ(p_{1})^{*}$
$*\displaystyle\simeq(p_{2})_{!}\circ(-\otimes((p_{1})^{*}X\otimes(p_{2})^{*}Y))\circ(p_{1})^{*}$
@@ 6.3. Bimodule perspective on the $(\infty,2)$-category of twisted spectra
$\mathbf{ThMod}$
$\operatorname{\mathbf{Pr}}_{\mathrm{St}}^{L}$
objects
$\mathbb{E}_{1}$-algebra $\operatorname{Th}(\alpha)$
presentable stable $\infty$-category
@@
1-morphism
$\operatorname{Th}(\alpha)-\operatorname{Th}(\beta)$-bimodule $M$
colimit preserving functor
@@
2-morphism
bimodule map $M\to N$
natural transformation
@@
identity 1-morphism
diagonal bimodule $\operatorname{Th}(\alpha)$
identity functor $1:\operatorname{Mod}_{\operatorname{Th}(\alpha)}\to\operatorname{Mod}_{\operatorname{Th}(\alpha)}$
vertical composition
vertical composition of
@@
horizontal composition on 1-morphism
composition of functors
tensoring over the joint common action

### v2

@@ 6.2. Dualizability of Twisted Spectra
$\Psi_{X}\circ\Phi_{Y}\simeq\Phi_{A_{!}(X\otimes Y)}\quad\text{and}\quad\Phi_{Y}\circ\Psi_{X}\simeq\Phi_{X\boxtimes Y}$
@@
$\Psi_{X}\circ\Phi_{Y}$ $\simeq A_{!}\circ(-\otimes Y)\circ(-\otimes X)\circ A^{*}$
$\simeq(-\otimes A_{!}(X\otimes Y))\circ A^{*}$
@@
$\Phi_{X\boxtimes Y}$ $=(p_{2})_{!}\circ(-\otimes(X\boxtimes Y))\circ(p_{1})^{*}$
$\simeq(p_{2})_{!}\circ(-\otimes((p_{1})^{*}X\otimes(p_{2})^{*}Y))\circ(p_{1})^{*}$
@@ 6.3. Bimodule perspective on the $(\infty,2)$-category of twisted spectra
$\mathbf{ThMod}$ $\operatorname{\mathbf{Pr}}_{\mathrm{St}}^{L}$
objects $\mathbb{E}_{1}$-algebra $\operatorname{Th}(\alpha)$ presentable stable $\infty$-category
@@
1-morphism $\operatorname{Th}(\alpha)-\operatorname{Th}(\beta)$-bimodule $M$ colimit preserving functor
@@
2-morphism bimodule map $M\to N$ natural transformation
@@
identity 1-morphism diagonal bimodule $\operatorname{Th}(\alpha)$ identity functor $1:\operatorname{Mod}_{\operatorname{Th}(\alpha)}\to\operatorname{Mod}_{\operatorname{Th}(\alpha)}$
vertical composition vertical composition of
@@
horizontal composition on 1-morphism composition of functors tensoring over the joint common action